# Pressure, Reference Points, and Risk-Taking Behavior: Evidence from Bench Press Competitions

### v1

@@ 2.3 Exogenous pressure
In the choice of attempt weight at the second attempt, first, each lifter knows the first attempt outcome and the second attempt weight of his one lower rank rival, which are uncorrelated with his unobserved characteristics such as his body condition at the competition date. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. Second, each lifter also knows the first attempt outcome of his one higher rank rival and expects the second attempt of the rival, which is uncorrelated with his unobserved characteristics. We assume that each lifter has a rational expectation of the higher rank rival’s attempt because it is predictable so that we treat the second actual attempt of the rival as the second actual attempt known to the lifter. As the first attempt weight of his one higher rank rival heavier than his current best weight, it is more difficult to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain.
@@
In the first attempt of lifting the attempt weight, first, each lifter knows the first attempt outcome and the second attempt weight of his one lower rank rival, which are uncorrelated with his unobserved characteristics such as his body condition at the competition date. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. If his second attempt fails, he is turned around the rival. Second, each lifter also knows the second attempt weight of his one higher rank rival, which is also uncorrelated with his unobserved characteristics. If his second attempt is successful, he has more chances to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain.
@@ 2.4 Data pattern
We illustrate data patterns of the choice of attempt weight and the success probability of the attempt weight conditional on the above exogenous pressure. For illustration, we show the case of the most popular competition category (Male, Raw, 24-39 age class). The data pattern is prevalent in the other competition categories, shown in the Appendix. We omit the data pattern for the first bench press attempts because the success probability at the stage is more than 75%. [^7] See Appendix for all plots for all competition categories for reference.
@@
Figure 2 displays success probabilities for second bench press attempts across different weight classes, showing the highest success rates when the attempted weight is slightly below the lifter’s personal best. Across almost all classes, success peaks around a distance of zero, with success probabilities typically ranging from 0.75 to 1.00 at this point. As the attempted weight deviates from the best, success rates drop significantly, often falling below 0.25 at more extreme deviations, particularly for heavier attempts. In heavier weight classes (120 kg and 120+ kg), the range of successful attempts is slightly broader, but the general trend of declining success with larger deviations remains consistent across all weight categories. While both pressured (“turning around” or “turned around”) or non-pressured groups follow similar patterns, the pressured group generally maintains slightly higher success probabilities, especially near the lifters’ personal best.
@@
Figure 3 illustrates more remarkable patterns for third bench press attempts. In comparison to the second attempt, the success rates in the third attempt are noticeably lower across all weight classes. While success rates during the second attempt peaked between 0.75 and 1.00 near the lifters’ personal best, the third attempt generally shows lower peak success rates, ranging between 0.5 and 0.8. Additionally, the drop in success rates as lifters deviate from their personal best is more pronounced in the third attempt, with success probabilities frequently falling below 0.25 at larger deviations. The distinction between the “turning around” groups (0 and 1) remains consistent, though the overall lower success rates make the difference less noticeable in the third attempt. Overall, lifters appear to struggle more in the third attempt, especially with larger deviations from their best, indicating a significant drop in success probability compared to the second attempt.
In summary, the above data patterns imply that when players are under pressure in some critical situation to change their own rank, they attempt more challenging weights and achieve them with higher probability.
@@
We examine the pressure exerted by rivals during competition. A natural question arises: who constitutes a rival for each lifter within a competition? Table 3 illustrates the distribution of rank changes for lifters between their second and third attempts. The majority of lifters maintain their rank, as shown by the high proportion of zero rank changes (68% in the second attempt and 82% in the third attempt). A smaller percentage of lifters experience a rank increase ($+1$) between attempts (15% in the second and 10% in the third), while the proportion of those gaining two or more ranks is notably lower (6% in the second and 2% in the third). Similarly, lifters dropping one rank ($-1$) account for 7% in the second attempt and 5% in the third, with even fewer dropping two or more ranks (4% and 2%, respectively). Based on this distribution, we focus on the immediate competitors, specifically the lifter ranked directly below and the one ranked directly above, as these rivals most significantly influence rank changes.
@@ 3 Theoretical prediction
We provide theoretical predictions based on expected utility and prospect theory. [^8] See the survey of O’Donoghue and Sprenger (2018) and DellaVigna (2018) for reference. The choice of attempt weight is analogous to a binary lottery choice from a set of lotteries ${(x,q(x))}$, where $x$ is the attempt weight and $q(x)\in[0,1]$ is the objective probability of success for the attempt, which is naturally assumed to be a decreasing function in $x$. The expected evaluation of lottery $(x,q(x))$ is
$U(x,q(x)|r)=v(x-r)q(x),$
where $v(\cdot)$ captures reference-dependent utility of outcome $x$ given reference point $r$. Following Tversky and Kahneman (1992), $v(\cdot)$ is specified as
$v(x-r)=\left\{\begin{array}[]{ll}(x-r)^{\alpha}&\text{if }x\geq r\\ -\lambda(r-x)^{\alpha}&\text{if }x<r.\end{array}\right.$
where $0<\alpha\leq 1$ and $\lambda>1$ capture diminishing sensitivity—concave curvature in the gain domain and a convex curvature in the loss domain—and loss aversion, respectively.
@@ 3.1 Lifting stage
Consider the attempt and lifting stages backward. We do not incorporate strategic decisions across stages to avoid complications making the situation intractable, but this would be an interesting research avenue. In the lifting stage, the reference point $r$ is the lifter’s current ranking, normalized to $r=0$, and a change in ranking induces $x\in\{-1,0,1\}$, so $\alpha$ does not matter in this setting. Let $p_{H}$ denote the ex-ante probability of the higher rival’s success, which is unknown to lifter $i$ in the lifting stage. Also, let $p_{L}$ denote the ex-ante probability of the lower rival’s success, which is unknown to lifter $i$ in the attempt stage.
Table 4 describes the utility tuple $(v_{1}(x),(v_{2}(x),v_{3}(x)))$ conditional on four scenarios determined by pressure from lower and higher rivals, where $v_{1}(x)$ is the utility when lifter $i$’s attempt weight $x$ fails, and $(v_{2}(x),v_{3}(x))$ is the utility when lifter $i$’s attempt is successful and the higher rival’s attempt either succeeds or fails. If there is no pressure from either lower or higher rivals (i.e., (No, No)), the current lifting attempt does not change lifter $i$’s rank whether successful or not, meaning $(v_{1}(x),(v_{2}(x),v_{3}(x)))=(0,(0,0))$. If there is pressure only from a lower rival (i.e., (Yes, No)), lifter $i$ experiences disutility $-\lambda(<-1)$ upon failure, but zero otherwise. If there is pressure only from a higher rival (i.e., (No, Yes)), lifter $i$ gains 1 if the higher rival fails, with the probability of $1-p_{H}$. If there is pressure from both lower and higher rivals, lifter $i$ experiences disutility $-\lambda(<-1)$ upon failure, gains 00 when both lifter $i$ and the higher rival succeed, and gains 1 if the higher rival fails, with the probability of $1-p_{H}$. Based on these scenarios, lifter $i$ has the strongest incentive to make an effort to avoid loss and gain when there is pressure from his lower and higher rivals.
@@ 3.2 Attempt stage
Next, consider the attempt stage. Let $x_{L}$ denote the lower rival’s actual attempt weight which is observable to lifter $i$ during the attempt stage. Also, let $x_{H}$ denote the higher rival’s expected attempt weight under lifter $i$’s rational expectations. Although this is not observable to lifter $i$’s attempt stage, we treat this as observable or highly predictable information to lifter $i$ [^9] Using our data, we discuss the predictability of higher rival’s attempt weight given observable information at lifter $i$’s attempt stage in Section 7. Additionally, let $y_{L}$, $y_{H}$, and $y_{i}$ represent the best outcome weights from earlier attempts within the competition for the lower rival, the higher rival, and lifter $i$, respectively, which are known in lifter $i$’s current attempt stage. The inequality conditions $y_{L}<y_{i}<y_{H}$, determined by rank order, and $y_{L}<x_{L}$, $y_{i}<x_{i}$, and $y_{H}<x_{H}$, meaning that the outcome weight must not exceed the corresponding attempt weight, must hold. Although we can assume that $x_{L}<x_{H}$, meaning that the lower rival does not attempt to overtake two higher-ranked rivals as inferred from our data, must hold, we keep generality and allow $x_{L}>x_{H}$. These conditions determine the joint probability of the outcome tuple as shown in Table 5.
Table 5: Joint probability of outcome tuple
Pressure $i$ $L$ $H$ Outcome $i$’ rank Probability
$x_{i}<\min\{x_{L},x_{H}\}$ S S S $x_{i}<\min\{x_{L},x_{H}\}$- $q(x)p_{L}p_{H}$
$x_{i}<\min\{x_{L},x_{H}\}$ S S F $x_{i}<\min\{x_{L},y_{H}\}$- $q(x)p_{L}(1-p_{H})P(x_{i}<y_{H})$
$x_{i}<\min\{x_{L},x_{H}\}$ S S F $y_{H}<x_{i}<x_{L}$ K $q(x)p_{L}(1-p_{H})P(x_{i}>y_{H})$
$x_{i}<\min\{x_{L},x_{H}\}$ S F S $y_{L}(<y_{i})<x_{i}<x_{H}$ K $q(x)(1-p_{L})p_{H}$
$x_{i}<\min\{x_{L},x_{H}\}$ S F F $y_{L}(<y_{i})<x_{i}<y_{H}$ K $q(x)(1-p_{L})(1-p_{H})P(x_{i}<y_{H})$
$x_{i}<\min\{x_{L},x_{H}\}$ S F F $y_{L}<y_{H}<x_{i}$ + $q(x)(1-p_{L})(1-p_{H})P(x_{i}>y_{H})$
$x_{i}<\min\{x_{L},x_{H}\}$ F S Any $y_{i}<\min\{x_{L},y_{H}\}$- $(1-q(x))p_{L}$
$x_{i}<\min\{x_{L},x_{H}\}$ F F Any $y_{L}<y_{i}<y_{H}$ K $(1-q(x))(1-p_{L})$
$x_{L}<x_{i}<x_{H}$ S S S $x_{L}<x_{i}<x_{H}$ K $q(x)p_{L}p_{H}$
$x_{L}<x_{i}<x_{H}$ S S F $x_{L}<x_{i}<y_{H}(<x_{H})$ K $q(x)p_{L}(1-p_{H})P(x_{i}<y_{H})$
$x_{L}<x_{i}<x_{H}$ S S F $\max\{x_{L},y_{H}\}<x_{i}$ + $q(x)p_{L}(1-p_{H})P(x_{i}>y_{H})$
$x_{L}<x_{i}<x_{H}$ S F S $y_{L}(<x_{L})<x_{i}<x_{H}$ K $q(x)(1-p_{L})p_{H}$
$x_{L}<x_{i}<x_{H}$ S F F $y_{L}<x_{i}<y_{H}(<x_{H})$ K $q(x)(1-p_{L})(1-p_{H})P(x_{i}<y_{H})$
$x_{L}<x_{i}<x_{H}$ S F F $y_{L}<y_{H}<x_{i}$ + $q(x)(1-p_{L})(1-p_{H})P(x_{i}>y_{H})$
$x_{L}<x_{i}<x_{H}$ F S S $y_{i}<x_{L}<x_{H}$- $(1-q(x))p_{L}p_{H}P(x_{L}>y_{i})$
$x_{L}<x_{i}<x_{H}$ F S S $x_{L}<y_{i}<x_{H}$ K $(1-q(x))p_{L}p_{H}P(x_{L}<y_{i})$
$x_{L}<x_{i}<x_{H}$ F S F $x_{L}<y_{i}<y_{H}$ K $(1-q(x))p_{L}(1-p_{H})P(x_{L}<y_{i})$
$x_{L}<x_{i}<x_{H}$ F S F $y_{i}<\min\{x_{L},y_{H}\}$- $(1-q(x))p_{L}(1-p_{H})P(x_{L}>y_{i})$
$x_{L}<x_{i}<x_{H}$ F F S $y_{L}<y_{i}<x_{H}$ K $(1-q(x))(1-p_{L})p_{H}$
$x_{L}<x_{i}<x_{H}$ F F F $y_{L}<y_{i}<y_{H}$ K $(1-q(x))(1-p_{L})(1-p_{H})$
$\max\{x_{L},x_{H}\}<x_{i}$ S Any Any $\max\{x_{L},x_{H}\}<x_{i}$ + $q(x)$
$\max\{x_{L},x_{H}\}<x_{i}$ F S Any $x_{L}<y_{i}<\min\{y_{H},x_{H}\}$ K $(1-q(x))p_{L}P(x_{L}<y_{i})$
$\max\{x_{L},x_{H}\}<x_{i}$ F S Any $y_{i}<\min\{x_{L},y_{H}\}$- $(1-q(x))p_{L}P(x_{L}>y_{i})$
$\max\{x_{L},x_{H}\}<x_{i}$ F F Any $y_{L}<y_{i}<\min\{y_{H},x_{H}\}$ K $(1-q(x))(1-p_{L})$
Notes: F = failed, S = success, K = keep. We ignore the equality case like $x_{L}=y_{i}$ in the probability because the body weight information works as a tie-breaker for the case.
Given the values of $x_{L},x_{H},y_{L},y_{H}$, and $y_{i}$ with utility tuple in Table 4, the expected evaluation of the lottery $(x_{i},q(x_{i}))$ for lifter $i$ with reference point $r_{i}$ is
$U(x_{i},q(x_{i})|r_{i})=v(x_{i}-r_{i})q(x_{i})$
=$=\left\{\begin{array}[]{ll}\Bigl{(}q(x_{i})p_{L}p_{H}+q(x_{i})p_{L}(1-p_{H})P(x_{i}<y_{H})+(1-q(x))p_{L}\Bigr{)}[-\lambda]&\text{if }x_{i}<x_{L},x_{H}\\ +\Bigl{(}q(x_{i})(1-p_{L})(1-p_{H})P(x_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(x_{i}))p_{L}p_{H}P(x_{L}>y_{i})+(1-q(x_{i}))p_{L}(1-p_{H})P(x_{L}>y_{i})\Bigr{)}[-\lambda]&\text{if }x_{L}<x_{i}<x_{H}\\ +\Bigl{(}q(x_{i})p_{L}(1-p_{H})P(x_{i}>y_{H})+q(x_{i})(1-p_{L})(1-p_{H})P(x_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(x_{i}))p_{L}P(x_{L}>y_{i})\Bigr{)}[-\lambda]+q(x_{i})[1]&\text{if}x_{L},x_{H}<x_{i}\end{array}\right.$ (6)
=$=\left\{\begin{array}[]{ll}\Bigl{(}q(x_{i})p_{L}p_{H}+q(x_{i})p_{L}(1-p_{H})P(x_{i}<y_{H})+(1-q(x))p_{L}\Bigr{)}[-\lambda]&\text{if }x_{i}<x_{L},x_{H}\\ +\Bigl{(}q(x_{i})(1-p_{L})(1-p_{H})P(x_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(x_{i}))p_{L}P(x_{L}>y_{i})\Bigr{)}[-\lambda]+\Bigl{(}q(x_{i})(1-p_{H})P(x_{i}>y_{H})\Bigr{)}[1]&\text{if }x_{L}<x_{i}<x_{H}\\ \Bigl{(}(1-q(x_{i}))p_{L}P(x_{L}>y_{i})\Bigr{)}[-\lambda]+q(x_{i})[1]&\text{if}x_{L},x_{H}<x_{i}\end{array}\right.$ (11)
Although the representation of the expected evaluation of the lottery in Equation (11) may appear complex, in particular, when $x_{L}<x_{i}<x_{H}$, it reveals that there are discontinuities at $x_{L}$ and $x_{H}$. Additionally, it is crucial to note that $P(x_{i}>y_{H})$ (that is, chance of the higher rival’s being turned around by the lifter) and $P(x_{L}>y_{i})$ (that is, chance of the lower rival’s turning around the lifter) are key factors that capture the pressure exerted by lower and higher rivals, respectively. [^10] Even without further specification, Equation (11) offers theoretical predictions regarding the optimal choice of attempt weight $x_{i}$ for each domain of $x_{i}$. First, if lifter $i$ aims to choose $x_{i}$ such that $\max\{x_{L},x_{H}\}<x_{i}$, the optimal choice for $x_{i}$ is $x_{H}+\kappa$, where $\kappa$ is the minimum weight increment as specified by the competition rules, given that $q(x_{i})$ is a decreasing function of $x_{i}$. Second, if lifter $i$ chooses $x_{i}$ such that $x_{i}<\min\{x_{L},x_{H}\}$, the optimal choice is the $x_{i}$ that maximizes $q(x_{i})P(x_{i}>y_{H})$, which leads to $x_{i}=y_{H}+\kappa$. Third, if lifter $i$ selects $x_{i}$ such that $x_{L}<x_{i}<x_{H}$, the optimal choice is $x_{i}$ that maximizes $q(x_{i})(1-p_{H})P(x_{i}>y_{H})$, which also results in $x_{i}=y_{H}+\kappa$. Further analysis to precisely determine the analytical optimal attempt weight $x_{i}$ across all domains of attempt weights would require additional specifications for $q(\cdot),p_{H},p_{L},\lambda$, and other parameters, which increases the complexity of the modeling, identification, and estimation processes, and is beyond the scope of this study and our data.

### v2

@@ 2.3 Exogenous pressure
In the choice of attempt weight at the second attempt, first, each lifter knows the first attempt outcome and the second attempt weight of his one lower rank rival. These are uncorrelated with the lifter’s unobservable characteristics, such as his body condition on the competition date, because the rival declares his second attempt weight before observing the outcome of the lifter’s second attempt, which—if observable—could serve as a proxy for the lifter’s condition. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. Second, each lifter expects the second attempt weight of his one higher rank rival, although this value is not yet observed. We assume that each lifter can make a prediction of a higher-ranked rival’s action based on information about the rival’s past attempts and a rich set of covariates. [^7] The assumption is supported from our data in Section 7.2. Strategic reasoning in intermediate weight selections is also inherently limited, as each subsequent attempt must exceed the current attempt. While incorporating dynamic and strategic considerations is theoretically appealing, doing so introduces significant complexity and falls beyond the scope of our empirical analysis. Because these predictions are based solely on observable characteristics, they are independent of the lifter’s own unobserved factors, such as temporary performance conditions. As the first attempt weight of his one higher rank rival is heavier than his current best weight, it is more difficult to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain.
@@
In the second attempt of lifting the attempt weight, each lifter knows the first and second attempt outcomes of his one lower rank rival. These are uncorrelated with the lifter’s unobserved characteristics, such as his body condition on the competition date, because the lower-ranked rival completes both attempts without observing the lifter’s second outcome. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. If his second attempt fails, he is turned around by the rival. Second, each lifter also knows the second attempt weight of his one higher rank rival. This declared weight is uncorrelated with the lifter’s unobserved characteristics, because at the time of declaration, the higher-ranked rival has not yet observed the lifter’s second outcome. If his second attempt is successful, he has more chances to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain.
@@ 2.4 Data pattern
We illustrate data patterns of the choice of attempt weight and the success probability of the attempt weight conditional on the above exogenous pressure. For illustration, we show the case of the most popular competition category (Male, Raw, 24-39 age class). The data pattern is prevalent in the other competition categories. We omit the data pattern for the first bench press attempts because the success probability at that stage is more than 75%.
@@
Figure 2 displays success probabilities for second bench press attempts across different weight classes, showing the highest success rates when the attempted weight is slightly below the lifter’s personal best. Across almost all classes, success peaks around a distance of zero, with success probabilities typically ranging from 0.75 to 1.00 at this point. As the attempted weight deviates from the best, success rates drop significantly, often falling below 0.25 at more extreme deviations, particularly for heavier attempts. In heavier weight classes (120 kg and 120+ kg), the range of successful attempts is slightly broader, but the general trend of declining success with larger deviations remains consistent across all weight categories. While both pressured (“turning around” or “turned around”) and non-pressured groups follow similar patterns, the pressured group generally maintains slightly higher success probabilities, particularly near the lifters’ personal best. However, variations exist across different weight categories, with some distances exhibiting no clear advantage for the pressured group.
@@
Figure 3 illustrates more distinct patterns for third bench press attempts. Compared to the second attempt, success rates in the third attempt are noticeably lower across all weight classes. While success rates during the second attempt peaked between 0.75 and 1.00 near the lifters’ personal best, the third attempt generally shows lower peak success rates, ranging between 0.5 and 0.8. Additionally, the decline in success rates as lifters deviate from their personal best is more pronounced in the third attempt, with success probabilities frequently dropping below 0.25 at larger deviations. The distinction between the “turning around” groups (0 and 1) remains visible, though the overall lower success rates make the difference less pronounced in the third attempt. Overall, lifters appear to struggle more in the third attempt, particularly when attempting weights significantly different from their personal best, indicating a notable drop in success probability compared to the second attempt.
In summary, the data suggests that when lifters face pressure in critical situations—such as attempting to improve their rank—they tend to choose more challenging weights. However, unlike the second attempt, where pressured lifters often maintain higher success rates, the third attempt shows a more substantial decline in overall success probabilities, suggesting increased difficulty in securing successful lifts under pressure.
@@
We examine the pressure exerted by rivals during competition. A natural question arises: who constitutes a rival for each lifter within a competition? Table 3 illustrates the distribution of rank changes for lifters between their second and third attempts. The majority of lifters maintain their rank, as shown by the high proportion of zero rank changes (67% in the second attempt and 72% in the third attempt). A smaller percentage of lifters experience a rank increase ($+1$) between attempts (14% in the second and 10% in the third), while the proportion of those gaining two or more ranks is notably lower (5% in the second and 3% in the third). Similarly, lifters dropping one rank ($-1$) account for 8% in the second attempt and 9% in the third, with even fewer dropping two or more ranks (5% and 6%, respectively). Based on this distribution, we focus on the immediate competitors, specifically the lifter ranked directly below and the one ranked directly above, as these rivals most significantly influence rank changes.
@@ 3 Theoretical prediction
In this section, we derive empirical predictions based on expected utility theory and prospect theory. [^8] See the survey of O’Donoghue and Sprenger (2018) and DellaVigna (2018) for reference. The choice of an attempt weight is analogous to a binary lottery choice drawn from a set of lotteries $\left(w,q(w)\right)$, where $w$ represents the attempt weight and $q(w)\in[0,1]$ denotes the probability of success. We assume that the success probability is decreasing in weight, i.e., $q^{\prime}(w)<0$. Following Tversky and Kahneman (1992), we adopt a rank-dependent value function $v(\cdot)$ with reference point $r$ defined as
$v(x)=\begin{cases}(x-r)^{\alpha}&\mbox{if\ }x\geq r\\ -\lambda(r-x)^{\alpha}&\mbox{if\ }x<r,\end{cases}$
where $0<\alpha\leq 1$ and $\lambda>1$ capture diminishing sensitivity—concave curvature in the gain domain and a convex curvature in the loss domain—and loss aversion, respectively. Here, $x$ denotes the rank of lifter $i$ at the end of the stage and $r$ is the intermediate rank of $i$ before this stage. In our application, the change in rank $(x-r)$ takes values in $\{-1,0,1\}$.
@@ 3.1 Lifting stage
Consider the attempt and lifting stages backward. We do not incorporate strategic decisions across stages to avoid complications making the situation intractable, but this would be an interesting research avenue. As noted earlier, the change in ranking takes values in $\{-1,0,1\}$, meaning that $\alpha$ does not affect this stage. Let $q(w_{H})$ denote the ex-ante probability that the higher-ranked rival successfully lifts the attempted weight $w_{H}$, which is unknown to lifter $i$ at the time of their own attempt. Similarly, let $q(w_{L})$ denote the ex-ante probability that the lower-ranked rival succeeds in lifting the attempted weight $w_{L}$, which is also unknown to lifter $i$ at the attempt stage.
@@
Table 4 describes the utility tuple $(v_{1}(x),(v_{2}(x),v_{3}(x)))$ conditional on four scenarios determined by pressure from lower and higher rivals, where $v_{1}(x)$ is the utility when lifter $i$’s attempt weight $x$ fails, and $(v_{2}(x),v_{3}(x))$ is the utility when lifter $i$’s attempt is successful and the higher rival’s attempt either succeeds or fails. For simplicity, we assume $\alpha=1$, which does not affect our prediction. If there is no pressure from either lower or higher rivals (i.e., (No, No)), the current lifting attempt does not change lifter $i$’s rank whether successful or not, meaning $(v_{1}(x),(v_{2}(x),v_{3}(x)))=(0,(0,0))$. If there is pressure only from a lower rival (i.e., (Yes, No)), lifter $i$ experiences disutility $-\lambda(<-1)$ upon failure, but zero otherwise. If there is pressure only from a higher rival (i.e., (No, Yes)), lifter $i$ gains 1 if the higher rival fails, with the probability of $1-q(w_{H})$. If there is pressure from both lower and higher rivals, lifter $i$ experiences disutility $-\lambda(<-1)$ upon failure, gains 00 when both lifter $i$ and the higher rival succeed, and gains 1 if the higher rival fails, with the probability of $1-q(w_{H})$. Based on these scenarios, lifter $i$ has the strongest incentive to make an effort to avoid loss and gain when there is pressure from his lower and higher rivals. Based on these scenarios, lifter $i$ has the strongest incentive to exert effort to avoid loss and seek gains when under pressure from both lower- and higher-ranked rivals. If $\lambda>1$, the incentive to exert effort to avoid loss is further amplified.
@@ 3.2 Attempt stage
We now examine the attempt stage, focusing on how lifter $i$’s choice of attempt weight is influenced by their competitors. For simplicity, we analyze the effects of the lower and higher rivals separately. This distinction is justified by the empirical analysis, where both sources of pressure are incorporated simultaneously into the model. By conditioning on the influence of one competitor while examining the effect of the other, we can isolate their respective impacts on the lifter’s decision-making. [^9] While it is possible to construct a model that simultaneously accounts for both rivals, doing so significantly increases the complexity of the framework. To maintain analytical tractability and improve interpretability, we present the separate analyses in the main text. A more comprehensive model incorporating both rivals jointly is provided in Appendix A.1.
3.2.1 Response to a lower rival
Let $w_{L}$ denote the attempt weight of the lower rival and let $y_{L}$ and $y_{i}$ be the best outcome weights from earlier attempts within the competition for the lower rival and lifter $i$, respectively. The inequality conditions $y_{L}<y_{i}$, determined by rank order, and $y_{L}\leq w_{L}$ and $y_{i}\leq w_{i}$, meaning that the outcome weight must not exceed the corresponding attempt weight, must hold, given by the current competition rule. We consider three cases: [^10] If $w_{i}=w_{L}$, the lifter with the lighter bodyweight receives the higher ranking. Such tie-weight situations are categorized into the three cases described above.
Case 1: $w_{L}>w_{i}\geq y_{i}$. Even if lifter $i$ succeeds, his attempt weight $w_{i}$ remains below the lower rival’s $w_{L}$; thus, his rank depends solely on the lower rival’s success probability $q(w_{L})$. The expected utility is
$EU|_{w_{L}>w_{i}\geq y_{i}}=$ $\left[1-q(w_{L})\right]\cdot 0+q(w_{L})\cdot[-\lambda]$
$=$ $-\lambda q(w_{L})\leq 0.$
Case 2: $w_{i}>w_{L}>y_{i}$. In this case, if lifter $i$ succeeds his rank remains unchanged; however, if he fails and the lower rival succeeds, he incurs a loss. Thus,
$EU|_{w_{i}>w_{L}>y_{i}}=$ $q(w_{i})\cdot 0+\left[1-q(w_{i})\right]\left[1-q(w_{L})\right]\cdot 0+\left[1-q(w_{i})\right]q(w_{L})\cdot[-\lambda]$
$=$ $-\lambda\left[1-q(w_{i})\right]q(w_{L})\leq 0.$
Case 3: $w_{i}\geq y_{i}>w_{L}$. Here, even if the lower rival succeeds, there is no change in rank; hence,
$EU|_{w_{i}\geq y_{i}>w_{L}}=0.$
Since $w_{L}$ is exogenously determined from the viewpoint of lifter $i$, we obtain the ordering
$EU\mid_{w_{i}\geq y_{i}>w_{L}}\geq EU\mid_{w_{i}>w_{L}>y_{i}}\geq EU\mid_{w_{L}>w_{i}\geq y_{i}}.$
Moreover, note that
$\frac{\partial EU\mid_{w_{i}>w_{L}>y_{i}}}{\partial w_{i}}=\lambda q(w_{L})q^{\prime}(w_{i})\leq 0,$
since $q^{\prime}(w_{i})<0$. Therefore, lifter $i$ will declare an attempt weight that is slightly heavier than that of the lower rival. Note that the value of $\lambda$ does not directly affect the level of the chosen attempt weight. As long as $\lambda>0$, the prediction remains informative and aligned with observed behavior.
3.2.2 Response to a higher rival
Let $w_{H}$ denote the declared weight of the higher rival. Although this is not observable to lifter $i$’s attempt stage, we treat this as highly predictable information to lifter $i$ [^11] Using our data, we discuss the predictability of higher rival’s attempt weight given observable information at lifter $i$’s attempt stage in Section 7. Additionally, let $y_{H}$ be the higher rival’s best outcome weights up to the previous stage. The following cases arise:
Case 1: $w_{H}\geq y_{H}>w_{i}$. When lifter $i$’s attempt weight $w_{i}$ is below the higher rival’s outcome up to the previous stage, there is no change in rank, so
$EU|_{w_{H}\geq y_{H}>w_{i}}=0.$
Case 2: $w_{H}>w_{i}>y_{H}$. In this case, if lifer $i$ succeeds and the higher rival fails, he overtakes the higher rival; otherwise, his rank remains unchanged. Hence,
$EU|_{w_{H}>w_{i}>y_{H}}=$ $q(w_{i})\left[1-q(w_{H})\right]\cdot 1$
$=$ $q(w_{i})\left[1-q(w_{H})\right]\geq 0.$
Case 3: $w_{i}>w_{H}\geq y_{H}$. By declaring a weight higher than the higher rival’s, lifter $i$ secures the rank improvement if he succeeds. Thus,
$EU|_{w_{i}>w_{H}\geq y_{H}}=$ $q(w_{i})\cdot 1+\left[1-q(w_{i})\right]\cdot 0$
$=$ $q(w_{i})\geq 0.$
Let $w_{i}^{+}$ denote a choice such that $w_{i}>w_{H}$ and $w_{i}^{-}$ denote a choice with $w_{i}<w_{H}$. Comparing the expected utilities, we have:
$EU|_{w_{i}>w_{H}\geq y_{H}}>EU|_{w_{H}>w_{i}>y_{H}}\Leftrightarrow\frac{q(w_{i}^{+})}{q(w_{i}^{-})}>1-q(w_{H}),$
noting that $q(w_{i}^{+})<q(w_{i}^{-})$ since higher attempt weight implies a lower success probability and therefore $q(w_{i}^{+})/q(w_{i}^{-})\in(0,1)$.
Given $w_{H}$ is fixed, if the higher rival’s success probability $q(w_{H})$ is sufficiently high and/or if the success probability $q(w_{i})$ does not change significantly whether lifter $i$ chooses a weight that exceeds $w_{H}$ or one that remains below $w_{H}$, then we obtain
$EU|_{w_{i}>w_{H}\geq y_{H}}\geq EU|_{w_{H}>w_{i}>y_{H}}\geq EU|_{w_{H}\geq y_{H}>w_{i}}.$
Otherwise, if $q(w_{i})$ declines steeply when $w_{i}$ exceeds $w_{H}$, then the reverse ordering holds:
$EU|_{w_{H}>w_{i}>y_{H}}\geq EU|_{w_{i}>w_{H}\geq y_{H}}\geq EU|_{w_{H}\geq y_{H}>w_{i}}.$
Furthermore, note that
$\frac{\partial EU|_{w_{i}>w_{H}\geq y_{H}}}{\partial w_{i}}=q^{\prime}(w_{i})<0,\ \text{and}\ \frac{\partial EU|_{w_{H}>w_{i}>y_{H}}}{\partial w_{i}}=q^{\prime}(w_{i})\left[1-q(w_{H})\right]\leq 0.$
Therefore, lifter $i$ will declare a weight that is slightly heavier than either the higher rival’s (expected) declared weight or the higher rival’s best outcome up to the previous stage (i.e., $y_{H}$), depending on the higher rival’s success probability $q(w_{H})$ and on the sensitivity of $q(w_{i})$ to changes in $w_{i}$.